# What MLLMs Learn about When they Learn about Multimodal Reasoning: Perception, Reasoning, or their Integration?

### v2

@@ Abstract
Multimodal reasoning models have recently shown promise on challenging domains such as olympiad-level geometry, yet their evaluation remains dominated by aggregate accuracy, a single score that obscures where and how models are improving. We introduce MathLens, a benchmark designed to disentangle the subskills of multimodal reasoning while preserving the complexity of textbook-style geometry problems. The benchmark separates performance into three components: Perception: extracting information from raw inputs, Reasoning: operating on available information, and Integration: selecting relevant perceptual evidence and applying it within reasoning. To support each test, we provide annotations: visual diagrams, textual descriptions to evaluate reasoning in isolation, controlled questions that require both modalities, and probes for fine-grained perceptual skills, all derived from symbolic specifications of the problems to ensure consistency and robustness. Our analysis reveals that different training approaches have uneven effects: First, reinforcement learning chiefly strengthens perception, especially when supported by textual supervision, while textual SFT indirectly improves perception through reflective reasoning. Second, reasoning improves only in tandem with perception. Third, integration remains the weakest capacity, with residual errors concentrated there once other skills advance. Finally, robustness diverges: RL improves consistency under diagram variation, whereas multimodal SFT reduces it through overfitting. We will release all data and experimental logs.
@@ 1 Introduction
Recent advances in reasoning with Large Language Models (LLMs) have yielded remarkable progress in challenging domains such as Olympiad-level mathematics (Mathematical Association of America (2025)), graduate-level scientific question answering (Rein et al. (2024)), and multi-step program synthesis (Austin et al. (2021); Chen et al. (2021)). Motivated by these successes, a natural extension is to adapt similar training paradigms to Multimodal Large Language Models (MLLMs), equipping them with reasoning capabilities over both text and visual inputs. Tasks such as mathematical problem solving (Lu et al. (2024); Wang et al. (2024b)) and visual puzzles (Dao & Vu (2025); Ghosal et al. (2024); Feng et al. (2025)) illustrate the potential of this direction, giving rise to methods that adapt Supervised FineTuning (SFT) (Sun et al. (2025); Chung et al. (2025)) and Reinforcement Learning (RL) (Deng et al. (2025); Meng et al. (2025)) for multimodal reasoning, including sequentially combining both stages for enhanced reasoning.
@@
Figure 1: MathLens decomposes multimodal reasoning errors into perception, reasoning, and integration, revealing capacity-specific shifts after fine-tuning that are hidden by aggregate accuracy. Each training strategy affects capacities differently; e.g., text SFT yields a minor gain ($\nearrow$) in reasoning but harms ($\downarrow$) integration (model details in appendix D).
@@
However, unlike LLMs where reasoning-oriented training consistently yields substantial gains, multimodal reasoning training exhibits highly variable outcomes. This motivates analysis of how different training strategies influence specific skills. To this end, multimodal reasoning may be decomposed into perception, reasoning, and their integration, as each corresponds to a distinct source of error. Existing benchmarks, however, rarely adopt such a decomposition and instead primarily report aggregate accuracy, which obscures these distinctions. Some vary input modalities to approximate skill-specific testing, but without strict controls they fail to isolate capacities and offer limited diagnostic value (see Appendix B). Consequently, it remains unclear when multimodal reasoning training benefits MLLMs and which training signals or architectural choices are responsible.
@@
To close this gap, we present MathLens, a controlled benchmark of geometry problems that isolates perception, reasoning, and their integration (fig. 1). Starting from the symbolic semantic state of (Zhang et al., 2023), we build four aligned annotations (section 2.2): (i) diagrams rendered from geometric constraints to test perception, (ii) a definitionally equivalent textual description to test reasoning when perception is trivialized, (iii) multimodal questions requiring both modalities, and (iv) fine-grained probes targeting recovery of visual details. To further test robustness against visual modifications, MathLens introduces semantic diagram modifications that alter visual form while preserving task correctness. By grounding the benchmark in geometry, MathLens retains authentic task complexity and enables rigorous comparison to prior benchmarks.
@@
MathLens demonstrates that training strategies shape multimodal reasoning capacities in distinct ways. 1) Perception is primarily boosted by reinforcement learning, with larger gains when strong textual SFT has already established reasoning competence (section 3.2), while textual SFT itself, despite lacking visual input, indirectly strengthens perception through reflective reasoning (section 3.3). 2) Reasoning improves in tandem with perception under RL, but does not exhibit distinct additional gains beyond those coupled improvements (section 3.4). 3) Integration remains the least improved capacity: RL offers little benefit, and as perception and reasoning advance together, residual errors increasingly concentrate in integration, leaving it as the dominant failure mode (section 3.4). 4) Robustness diverges across strategies, with RL enhancing consistency under diagram variation, whereas multimodal SFT reduces robustness through overfitting (section 3.5).
Our contributions are:
@@
  • Framework for multi-axial evaluation of multimodal reasoning: A new benchmark that disentangles perception, reasoning, and integration, enabling analysis beyond aggregate accuracy.
@@
  • MathLens dataset: A public release of all problems, annotations, evaluation scripts, and perturbation generators, establishing a reproducible resource for the community.
@@ 2 MathLens
Most multimodal reasoning benchmarks report only aggregate accuracy, obscuring whether errors stem from perception (extracting information from inputs), reasoning (operating on extracted information), or their interaction. Our benchmark is designed to disentangle these subskills while preserving realistic problem contexts. MathLens comprises 926 geometry problems, each presented with eight visual modifications and an average of $\sim 7.03$ visual probes per problem.
[…]
@@ 2.1 Dataset formalization
with $\varphi_{k}=$ (compute $\angle A$) and $S_{k}^{\text{q}}=\{s_{k,1}\}$, the question can be $\angle ABC=50^{\circ}$, what is $\angle A$?” The ground-truth answer is defined by applying the operator to the full semantic state,
@@
so each probe targets a single atom of $S_{k}$ and its gold answer is the atom’s valuation. For example, $q^{\mathrm{perc}}_{k,1}=$ “Which angle has measure of 50? A. ABC B. ABD” with $a^{\mathrm{perc}}_{k,1}=$ A. These probes directly test whether a model has recovered the components of $S_{k}$ from the observed context.
@@
Having access to the semantic state $S_{k}$ allows us to design tests that separate perception, reasoning, and their integration, as shown in Figure 2.
Perception. Probe questions $Q^{\mathrm{perc}}k$ with gold answers $a^{\mathrm{perc}}{k,i}$ are derived from the atomic facts listed in $S_{k}$. These probes check if the model can recover the specific facts in $S_{k}$ that are needed to solve the problem from the given input. Errors here indicate pure perceptual failures.
@@
Figure 3: Sample data generation process in MathLens. From a semantic state representation, we build controlled text descriptions, perception probes, and questions with no overlap with visual content. Also, new diagrams are rendered from the semantic state to avoid visual familiarity effects [^2] In this example, an obtuse angle is drawn acute. Such non-canonical diagrams block visual estimation and enforce geometric deduction, consistent with Olympiad practice (“not drawn to scale,” e.g., AMC)..
Integration. We isolate integration effects by combining accuracy on $(C^{\mathrm{img}}_{k},Q_{k})$ with auxiliary perception and reasoning measures. Conditional on success in perception probes and text-only reasoning, any remaining errors on the full task are treated as integration errors; this defines integration as the residual after controlling for perception and reasoning under standard coverage assumptions.
@@
Second, the symbolic representation $S_{k}$ allows us to apply semantic perturbations of the context. Let $\tau\in\mathcal{T}_{\mathrm{AP}}$ denote a transformation from the set of admissible perturbations (e.g., relabeling points, rotating a diagram, or permuting equivalent elements). Applying $\tau$ yields a perturbed specification $S^{\prime}_{k}=\tau(S_{k})$, while preserving the problem semantics so that
@@ 2.2 Data Generation Pipeline
Thus, we build on latent semantics $S_{k}$ for a set of practical multimodal reasoning problems, matching problem domain and complexity with popular benchmarks in the literature (Zhang et al. (2024); Lu et al. (2024)). Each instance is first specified symbolically as $S_{k}$ and $\varphi_{k}$, then rendered into its observable forms: diagrams $C^{\mathrm{img}}_{k}$, textual descriptions $C^{\mathrm{txt}}_{k}$, and multiple types of questions.
Data source (fig. 4 a).
We build on FormalGeo-7K (Zhang et al. (2023)), which provides symbolic annotations for geometry problems. Each diagram cue or condition is encoded as a predicate (e.g., $\textsc{Collinear}(AB,BC)$, $\textsc{EqualLength}(AB,CD)$), forming the semantic state $S_{k}$. This yields realistic problems with formal representations from which we build the required artifacts.
Questions (fig. 4 b).
@@
To test robustness against distractions, we alter the semantic state $S_{k}$ to generate diagram variants. Modifications include adding auxiliary geometric elements, applying flips or rotations, merging instances, and relabeling points, while preserving the ground-truth answer. All variants are manually screened to discard visually invalid cases.
2.3 MathLens-General: A complementary general-domain set
The main dataset, MathLens, solely consists of geometry problems. While this symbolic structure enables precise control, it also risks overfitting analyses to a narrow domain. To mitigate this limitation, we construct MathLens-General, a set of multimodal reasoning problems spanning diverse domains. Unlike MathLens, it cannot maintain the same rigor afforded by formal semantic states, but is instead curated from prior sources through a rigorous pipeline to ensure reliability and diversity. Curation procedures and experimental results are provided in Appendices C and E.
Figure 4: Impact of multimodal reasoning training on MathLens performance. We evaluate pretrained backbones alongside models finetuned for multimodal reasoning tasks, reporting accuracy (%). MathLens is sensitive to gains from multimodal reasoning–oriented finetuning.
@@ 3 Experiments
We focus on open-weight multimodal reasoning models in the 7–9B parameter range, along with the backbone counterparts. This setup is motivated by two factors: transparency of training methods and data, which is available only for open models, and the fact that most publicly released multimodal reasoning models fall within this size range. We assemble seven model families, yielding 13 checkpoints in total by including both SFT-only and RL-finetuned variants where available. We include Qwen-2.5-VL (Bai et al., 2025) and GLM-4.1V-Base (Hong et al., 2025) as backbone models. Direct multimodal Reinforcement Learning (RL) models comprise VL-Rethinker (Wang et al., 2025) and ShareVL-R1 (Yao et al., 2025). Models trained with multimodal Supervised Finetuning (SFT) followed by multimodal RL include Vision-R1 (Huang et al., 2025) and R1-Onevision (Yang et al., 2025), while those using textual SFT followed by multimodal RL include Revisual-R1 (Chen et al., 2025) [^3] Revisual-R1 adds additional textual RL, yet we group it as textual SFT $\rightarrow$ multimodal RL for consistency. and Open-Vision-Reasoner (OVR) (Wei et al., 2025). Finally, GLM-4.1V-Thinking (Hong et al., 2025) falls outside these categories, since its training data is not publicly disclosed.
We additionally evaluate six larger models, yielding eight runs in total by including both Gemini-2.5-Flash (Comanici et al., 2025) and Claude 4 Sonnet (Anthropic, 2025) in “thinking” and “non-thinking” modes: Qwen-2.5-VL-72B, VL-Rethinker-72B, Gemini-2.5-Flash, Claude 4 Sonnet, and GPT-O3/4O (OpenAI, 2025). For each model, we follow the recommended decoding configurations to generate both reasoning and final answers. Most models employ greedy decoding and yield deterministic outputs. Full experimental details and results are in Section E.4.
@@
We take it as baseline that current multimodal SFT models show limited reasoning following Chen et al. (2025): their data are easier than text-only sets, leading to weaker performance on reasoning benchmarks. Building stronger data is hindered by the lack of open multimodal reasoning traces (see Section E.1 for discussion).
3.1 Validating MathLens as a multimodal reasoning benchmark
Before using MathLens to study finetuning effects, we first validate it as a multimodal reasoning benchmark, showing that it captures the same skills as established benchmarks.

### v3

@@ Abstract
Evaluation of multimodal reasoning models is typically reduced to a single accuracy score, implicitly treating reasoning as a unitary capability. We introduce MathLens, a benchmark of textbook-style geometry problems that exposes this assumption by operationally decomposing performance into Perception, Reasoning, and Integration. Each problem is derived from a symbolic specification and accompanied by visual diagrams, text-only variants, multimodal questions, and targeted perceptual probes, enabling controlled measurement of each component. Using this decomposition, we show that common training strategies induce systematically different capability profiles that are invisible under aggregate accuracy. Reinforcement learning primarily improves perceptual grounding and robustness to diagram variation, while textual SFT yields gains through reflective reasoning. In contrast, as perception and reasoning improve, a growing fraction of remaining errors fall outside these components and are categorized as integration. These results suggest that apparent progress in multimodal reasoning reflects shifting balances among subskills rather than uniform advancement, motivating evaluation beyond scalar accuracy.
Large Language Models,Multimodal Reasoning,Mathematical Reasoning,Compositionality
@@ 1 Introduction
Recent advances in reasoning with Large Language Models (LLMs) have yielded remarkable progress in challenging domains such as Olympiad-level mathematics ((Mathematical Association of America, 2025)), graduate-level scientific question answering ((Rein et al., 2024)), and multi-step program synthesis ((Austin et al., 2021; Chen et al., 2021)). Motivated by these successes, a natural extension is to adapt similar training paradigms to Multimodal Large Language Models (MLLMs), equipping them with reasoning capabilities over both text and visual inputs. Tasks such as mathematical problem solving ((Lu et al., 2024; Wang et al., 2024b)) and visual puzzles ((Dao & Vu, 2025; Ghosal et al., 2024; Feng et al., 2025)) illustrate the potential of this direction, giving rise to methods that adapt Supervised FineTuning (SFT) ((Sun et al., 2025; Chung et al., 2025)) and Reinforcement Learning (RL) ((Deng et al., 2025; Meng et al., 2025)) for multimodal reasoning, including sequentially combining both stages for enhanced reasoning.
@@
Figure 1: MathLens decomposes multimodal reasoning errors into perception, reasoning, and integration, revealing capacity-specific shifts after fine-tuning that are hidden by aggregate accuracy. Each training strategy affects capacities differently; e.g., text SFT yields a minor gain ($\nearrow$) in reasoning but harms ($\downarrow$) integration (model details in Appendix˜D).
@@
However, unlike LLMs where reasoning-oriented training yields consistent gains, multimodal reasoning training exhibits highly variable outcomes. This motivates analysis of how different training strategies influence specific skills. To this end, multimodal reasoning can be analyzed in terms of perception, reasoning, and their integration, each corresponding to a principal source of error. Existing benchmarks, however, rarely adopt such a decomposition and instead primarily report aggregate accuracy, which obscures the distinctions. Some vary input modalities to approximate skill-specific testing, but without strict controls they fail to isolate capacities and offer limited diagnostic value (see Appendix˜B). Consequently, it remains unclear which training strategies benefit multimodal reasoning and why.
A central challenge in addressing this question is the lack of domains where such a decomposition can be carried out with explicit semantic control. Geometry provides a natural setting for this analysis: its symbolic structure makes it possible to precisely specify perceptual facts, reasoning operators, and their interaction, and it already underpins most existing multimodal reasoning benchmarks (Lu et al., 2024; Wang et al., 2024b; Zhang et al., 2024). This level of control is essential for analyzing how training strategies differentially affect perception, reasoning, and their interaction. However, existing benchmarks do not fully exploit this structure to provide controlled, component-wise evaluation.
@@
To close this gap, we present MathLens, a controlled benchmark of geometry problems that separately probes perception, reasoning, and their integration (Figure˜1). Starting from the symbolic semantic state of (Zhang et al., 2023), we build four aligned annotations (Section˜2.2): (i) diagrams rendered from geometric constraints to primarily test perception, (ii) a definitionally equivalent textual description to test reasoning under trivialized perception, (iii) multimodal questions requiring both modalities, and (iv) fine-grained probes targeting recovery of visual details. To further test robustness against visual modifications, MathLens introduces semantic diagram modifications that alter visual form while preserving task correctness. By grounding the benchmark in geometry, MathLens retains authentic task complexity and enables rigorous comparison to prior benchmarks.
@@
MathLens demonstrates that training strategies shape multimodal reasoning capacities in distinct ways. 1) Perception is primarily boosted by reinforcement learning, with larger gains when strong textual SFT has already established reasoning competence (Section˜3.2), while textual SFT itself, despite lacking visual input, indirectly strengthens perception through reflective reasoning (Section˜3.3). 2) Reasoning improves in tandem with perception under RL, but does not exhibit distinct additional gains beyond those coupled improvements (Section˜3.4). 3) Integration remains the least improved capacity: RL offers little benefit, and as perception and reasoning improve, errors increasingly fall outside the scope of the measured perception and reasoning components, and are therefore categorized as integration. (Section˜3.4). 4) Robustness diverges across strategies, with RL enhancing consistency under diagram variation, whereas multimodal SFT reduces robustness through overfitting (Section˜3.5).
Our contributions:
@@
  • Framework for multi-axial evaluation of multimodal reasoning: A new benchmark that separately probes perception, reasoning, and integration, enabling analysis beyond aggregate accuracy.
@@
  • MathLens dataset: A curated benchmark with problems, annotations, and perturbation generators.
[…]
@@ 2 MathLens
Most multimodal reasoning benchmarks report only aggregate accuracy, obscuring whether errors stem from perception (extracting information from inputs), reasoning (operating on extracted information), or their interaction. Our benchmark is designed to separately probe these subskills while preserving realistic problem contexts. MathLens comprises 926 geometry problems, each presented with eight visual modifications and an average of $\sim 7.03$ visual probes per problem.
@@ 2.1 Dataset formalization
with $\varphi_{k}=$ (compute $\angle A$) and $S_{k}^{\text{q}}=\{s_{k,1}\}$, the question can be $\angle ABC=50^{\circ}$, what is $\angle A$?" The ground-truth answer is defined by applying the operator to the full semantic state,
@@
so each probe targets a single atom of $S_{k}$ and its gold answer is the atom’s valuation. For example, $q^{\mathrm{perc}}_{k,1}=$ “Which angle has measure of 50? A. ABC B. ABD" with $a^{\mathrm{perc}}_{k,1}=$ A. These probes directly test whether a model has recovered the components of $S_{k}$ from the observed context.
@@
Having access to the semantic state $S_{k}$ allows us to design tests that separately probe perception, reasoning, and their integration (Figure˜2).
Perception. Probe questions $Q^{\mathrm{perc}}k$ with gold answers $a^{\mathrm{perc}}{k,i}$ are derived from the atomic facts listed in $S_{k}$. These probes check if the model can recover the specific facts in $S_{k}$ that are needed to solve the problem from the given input. Errors here indicate perceptual failures.
@@
Integration. We isolate integration effects by combining accuracy on $(C^{\mathrm{img}}_{k},Q_{k})$ with auxiliary perception and reasoning measures. Conditional on success in perception probes and text-only reasoning, any remaining errors on the full task are treated as integration errors; this defines integration as a residual category after controlling for the measured perception and reasoning components under standard, non-exhaustive coverage assumptions.
Accordingly, integration should be interpreted as a conditional residual rather than a fully disentangled construct: it captures failures not explained by the measured perception probes and text-only reasoning, but may still subsume unmeasured perceptual factors.
@@
Second, the symbolic representation $S_{k}$ allows us to apply semantic perturbations of the context. Let $\tau\in\mathcal{T}_{\mathrm{AP}}$ denote a transformation from the set of admissible perturbations (e.g., relabeling points or rotating a diagram). Applying $\tau$ yields a perturbed specification $S^{\prime}_{k}=\tau(S_{k})$, while preserving the problem semantics so that
@@
Figure 3: Sample data generation process in MathLens. From a semantic state representation, we build controlled text descriptions, perception probes, and questions with no overlap with visual content. Also, new diagrams are rendered from the semantic state to avoid visual familiarity effects [^2] Here, an obtuse angle is drawn acute to prevent visual estimation and enforce geometric deduction, consistent with Olympiad practice (e.g., “not drawn to scale” in AMC)..
@@ 2.2 Data generation pipeline
Thus, we build on latent semantics $S_{k}$ for a set of practical multimodal reasoning problems, matching problem domain and complexity with popular benchmarks in the literature ((Zhang et al., 2024; Lu et al., 2024)). Each instance is first specified symbolically as $S_{k}$ and $\varphi_{k}$, then rendered into its observable forms: diagrams $C^{\mathrm{img}}_{k}$, textual descriptions $C^{\mathrm{txt}}_{k}$, and multiple types of questions.
Data source (Figure˜4 a).
We build on FormalGeo-7K ((Zhang et al., 2023)), which provides symbolic annotations for geometry problems. Each diagram cue or condition is encoded as a predicate (e.g., $\textsc{Collinear}(AB,BC)$, $\textsc{EqualLength}(AB,CD)$), forming the semantic state $S_{k}$. This yields realistic problems with formal representations from which we build the required artifacts.
Questions (Figure˜4 b).
@@
To test robustness to distractions, we generate diagram variants by altering the semantic state $S_{k}$, including adding auxiliary elements, applying flips or rotations, merging instances, and relabeling points, while preserving the ground-truth answer. All variants are manually screened for visual validity.
2.3 MathLens-NS: A non-symbolic set
The main dataset, MathLens, focuses on geometry problems, a domain selected intentionally for its well-defined symbolic structure and precise experimental controllability. This choice is methodological: geometry provides a natural and practical setting for isolating and analyzing the multimodal reasoning mechanisms studied in this work.
In addition, we construct MathLens-NS, a supplementary collection of multimodal reasoning problems drawn from diverse domains. Relative to MathLens, this dataset necessarily trades formal semantic rigor for broader domain inclusion, reflecting the intrinsic differences between geometry and less structured multimodal reasoning settings. As a result, MathLens-NS does not admit explicit semantic state specifications; instead, it is curated from prior sources through a consistent and rigorous filtering pipeline to ensure reliability and diversity. Curation procedures and experimental results are provided in Appendices˜C and E.
Figure 4: Impact of multimodal reasoning training on MathLens. We evaluate pretrained backbones alongside models finetuned for multimodal reasoning, reporting accuracy (%). MathLens is sensitive to gains from multimodal reasoning-oriented finetuning.
@@ 3 Experiments
We study open-weight multimodal reasoning models in the 7-9B range and their backbone counterparts, motivated by the transparency of open models and the fact that most released multimodal reasoners fall within this scale. We evaluate seven model families, totaling 13 checkpoints, including both SFT-only and RL-finetuned variants where available. We include Qwen-2.5-VL (Bai et al., 2025) and GLM-4.1V-Base (Hong et al., 2025) as backbone models. Direct multimodal Reinforcement Learning (RL) models comprise VL-Rethinker (Wang et al., 2025) and ShareVL-R1 (Yao et al., 2025). Models trained with multimodal Supervised Finetuning (SFT) followed by multimodal RL include Vision-R1 (Huang et al., 2025) and R1-Onevision (Yang et al., 2025), while those using textual SFT followed by multimodal RL include Revisual-R1 (Chen et al., 2025) [^3] Revisual-R1 adds additional textual RL, yet we group it as textual SFT $\rightarrow$ multimodal RL for consistency. and Open-Vision-Reasoner (OVR) (Wei et al., 2025). Finally, GLM-4.1V-Thinking (Hong et al., 2025) falls outside these categories because its training data is undisclosed.
We additionally evaluate six larger models, yielding eight runs in total by including both Gemini-2.5-Flash (Comanici et al., 2025) and Claude 4 Sonnet (Anthropic, 2025) in “thinking” and “non-thinking” modes: Qwen-2.5-VL-72B, VL-Rethinker-72B, Gemini-2.5-Flash, Claude 4 Sonnet, and GPT-O3/4O (OpenAI, 2025). For each model, we follow the recommended decoding configurations to generate both reasoning and final answers. Most models employ greedy decoding and yield deterministic outputs. Full experimental details and results are in Section˜E.3.
Figure 5: Correlation of MathLens with popular benchmarks. MathLens shows high correlation with standard multimodal reasoning benchmarks.
@@
We take it as baseline that current multimodal SFT models show limited reasoning following (Chen et al., 2025): their data are easier than text-only sets, leading to weaker performance on reasoning benchmarks. Building stronger data is hindered by the lack of open multimodal reasoning traces (see Section˜E.1 for discussion).
Statistical significance.
Because our primary claims concern changes induced by finetuning, we report paired bootstrap confidence intervals for all pre-post finetuning deltas over problem instances (Figure˜6, 7, and 10). Unless otherwise noted, all evaluations use deterministic decoding; accordingly, reported uncertainty reflects benchmark sampling variability rather than decoding stochasticity. Raw accuracies and diagnostic visualizations (Figure˜4, 9, and 13) are reported descriptively and are not subjected to statistical inference.
3.1 Validating MathLens
Before analyzing finetuning effects, we validate MathLens as a multimodal reasoning benchmark by showing it captures the same skills as established benchmarks.